# appReferences

### v3

@@ III-A Problem Setting
In this work, we focus on example-based control (EBC), a more difficult form of imitation learning where we have no access to rewards or demonstrations, but instead are given a set of example states of a completed task: $s^{*}\in\mathcal{B}^{*}$, where $\mathcal{B}^{*}\subseteq\mathcal{S}$ and $\lvert\mathcal{B}^{*}\rvert<\infty$. The goal is to (i) leverage $\mathcal{B}^{*}$ and $\mathcal{B}$ to learn or define a state-conditional reward function $\hat{R}:\mathcal{S}\to\mathbb{R}$ that satisfies $\hat{R}(s^{*})\geq\hat{R}(s)$ for all $(s^{*},s)\in\mathcal{B}^{*}\times\mathcal{B}$, and (ii) learn a policy $\hat{\pi}$ that maximizes the expected return $\hat{\pi}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\hat{R}(s_{t})\right]$.
For any policy $\pi$, we can define the value function and Q-function respectively to be $V^{\pi}(s)=\mathbb{E}_{\pi}\left[Q^{\pi}(s,a)\right]$ and $Q^{\pi}(s,a)=\hat{R}(s)+\gamma\mathbb{E}_{\mathcal{P}}\left[V^{\pi}(s^{\prime})\right]$, corresponding to the return-to-go from state $s$ (and action $a$). Temporal difference (TD) algorithms aim to estimate $V^{\pi}$ or $Q^{\pi}$ to evaluate a policy [27]. Given a reward model $\hat{R}(\cdot)$, we can say $\hat{R}(s^{*}),s^{*}\in\mathcal{B}^{*}$, indicates reward for successful states and $\hat{R}(s),s\in\mathcal{B}$, for all other states. Assuming that $s^{*}$ transitions to itself, then for policy evaluation with mean-squared error (MSE), we can write the TD targets for non-successful and successful states, $y:\mathcal{S}\times\mathcal{S}\to\mathbb{R}$, of Q-updates as
@@ IV-A1 Environments
We conduct experiments in a large variety of tasks and environments, including those originally used in LfGP [4] and RCE [1]. Specifically, the tasks in [4] involve a simulated Franka Emika Panda arm, a blue and green block, a fixed “bring” area for each block, and a small slot with $<$1 mm tolerance for inserting each block. This environment provides various manipulation tasks that share the same state-action space. The tasks in [1] are a modified subset of those from [9], involving a simulated Sawyer arm, and three of the Adroit hand tasks originally presented in [30]. We also generate three modified delta-position (dp) Adroit hand environments, because we found that policies learned in the original absolute-position environments lack finesse and exploit simulator bugs. Finally, we study drawer and door opening tasks with a real Franka Emika Panda.
@@ IV-C1 Q-Value Overestimation and Value Penalization
To verify that VP enforces $y(s,s^{\prime})\leq y(s^{*},s^{*})$, we took snapshots of each learned Unstack-Stack and Insert agent, for both VPACE and ACE, at 300k steps and ran each policy for a single episode, recording per-timestep Q-values. Instead of showing Q-values directly, in Figs. 8 and 9, we show $Q(s_{t},a_{t})-\mathbb{E}_{s^{*}\sim\mathcal{B}^{*}}\left[V(s^{*})\right]$, which should be at most 00 for $y(s,s^{\prime})\leq y(s^{*},s^{*})$ to hold. ACE clearly violates $y(s,s^{\prime})\leq y(s^{*},s^{*})$, while VPACE does not. Furthermore, Fig. 9 shows examples of the consequences of overestimated Q-values: for $(s_{t},a_{t})$ pairs that appear to be out-of-distribution (OOD), $Q(s_{t},a_{t})-\mathbb{E}_{s^{*}\sim\mathcal{B}^{*}}\left[V(s^{*})\right]>0$, and the resulting policy reaches these states instead of the true goal.
@@ A-B Discriminator-Free Reward Labels with Mean Squared Error TD Updates
The resulting reward function is the expected label $\hat{R}(s,a)=\mathbb{E}_{\hat{\mathcal{B}}\sim\text{Categorical}_{s}(\{\mathcal{B},\mathcal{B}^{*}\})}\left[\boldsymbol{1}(\hat{\mathcal{B}}=\mathcal{B}^{*})\right]$, where $\text{Categorical}_{s}$ is a categorical distribution corresponding to the probability that $s$ belongs to buffers $\mathcal{B},\mathcal{B}^{*}$. Consequently, only successful states $s^{*}$ yields positive reward and the corresponding optimal policy will aim to reach a successful state as soon as possible. If we further assume that $s^{*}$ transitions to itself, then for policy evaluation with mean-squared error (MSE), we can write the temporal difference (TD) target, $y$, of Q-updates with:
@@ B-F Additional Implementation Details
Target Entropy $-\text{dim}(a)$
@@
SQIL Labels (Section B-F $(-1,1)$
@@
DAC reward function: for VPACE-DAC and VP-DAC, although there are many options for reward functions that map $D$ to $\hat{R}$citeappghasemipourDivergenceMinimizationPerspective2019, following [4, 12, 15], we set the reward to $\hat{R}_{\mathcal{T}}(s)=\log\left(D_{\mathcal{T}}(s)\right)-\log\left(1-D_{\mathcal{T}}(s)\right)$.
@@
Reward Scaling of .1: we use a reward scaling parameter of .1 for all implementations. Coupled with a discount rate $\gamma=0.99$ (common for much work in RL), this sets the expected minimum and maximum Q values for SQIL to .$\frac{-.1}{1-\gamma}=-10$ and .$\frac{.1}{1-\gamma}=10$.
@@ B-F1 Maintaining $Q^{\pi}_{\text{max}}$, $Q^{\pi}_{\text{min}}$ Estimates for Value Penalization
Our approach to value penalization requires maintaining estimates for or choosing $Q^{\pi}_{\text{max}}$ and $Q^{\pi}_{\text{min}}$. In both DAC and SQIL, the estimate of $Q^{\pi}_{\text{max}}$ comes from taking the mini-batch of data from $\mathcal{B}^{*}$, passing it through the Q function, taking the mean, and then using a median moving average filter to maintain an estimate. The “$Q^{\pi}_{\text{max}}$, $Q^{\pi}_{\text{min}}$ num. filter points” value from Table IV refers to the size of this filter. We chose 50 and used it for all of our experiments. We set $Q^{\pi}_{\text{min}}$ to
@@
where in SQIL, $\min(\hat{R}(s))=\hat{R}(s)$ is set to 0 or -1, and in DAC, we maintain an estimate of the minimum learned reward $\min(\hat{R})$ using a median moving average filter with the same length as the one used for $Q^{\pi}_{\text{max}}$

### v4

@@ III-A Problem Setting
In this work, we focus on example-based control (EBC), a more difficult form of imitation learning where we have no access to rewards or demonstrations, but instead are given a set of example states of a completed task: $s^{*}\in\mathcal{B}^{*}$, where $\mathcal{B}^{*}\subseteq\mathcal{S}$ and $\lvert\mathcal{B}^{*}\rvert$$\infty$. The goal is to (i) leverage $\mathcal{B}^{*}$ and $\mathcal{B}$ to learn or define a state-conditional reward function $\hat{R}:\mathcal{S}\to\mathbb{R}$ that satisfies $\hat{R}(s^{*})\geq\hat{R}(s)$ for all $(s^{*},s)\in\mathcal{B}^{*}\times\mathcal{B}$, and (ii) learn a policy $\hat{\pi}$ that maximizes the expected return $\hat{\pi}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\hat{R}(s_{t})\right]$.
[…]
@@ IV-A1 Environments
We conduct experiments in a large variety of tasks and environments, including those originally used in LfGP [4] and RCE [1]. Specifically, the tasks in [4] involve a simulated Franka Emika Panda arm, a blue and green block, a fixed “bring” area for each block, and a small slot with <<1 mm tolerance for inserting each block. This environment provides various manipulation tasks that share the same state-action space. The tasks in [1] are a modified subset of those from [9], involving a simulated Sawyer arm, and three of the Adroit hand tasks originally presented in [30]. We also generate three modified delta-position (dp) Adroit hand environments, because we found that policies learned in the original absolute-position environments lack finesse and exploit simulator bugs. Finally, we study drawer and door opening tasks with a real Franka Emika Panda.
@@ IV-C1 Q-Value Overestimation and Value Penalization
To verify that VP enforces $y(s,s^{\prime})\leq y(s^{*},s^{*})$, we took snapshots of each learned Unstack-Stack and Insert agent, for both VPACE and ACE, at 300k steps and ran each policy for a single episode, recording per-timestep Q-values. Instead of showing Q-values directly, in Figs. 8 and 9, we show $Q(s_{t},a_{t})-\mathbb{E}_{s^{*}\sim\mathcal{B}^{*}}\left[V(s^{*})\right]$, which should be at most 0 for $y(s,s^{\prime})\leq y(s^{*},s^{*})$ to hold. ACE clearly violates $y(s,s^{\prime})\leq y(s^{*},s^{*})$, while VPACE does not. Furthermore, Fig. 9 shows examples of the consequences of overestimated Q-values: for $(s_{t},a_{t})$ pairs that appear to be out-of-distribution (OOD), $Q(s_{t},a_{t})-\mathbb{E}_{s^{*}\sim\mathcal{B}^{*}}\left[V(s^{*})\right]>0$, and the resulting policy reaches these states instead of the true goal.
@@ A-B Discriminator-Free Reward Labels with Mean Squared Error TD Updates
[…]
@@ B-F Additional Implementation Details
Target Entropy $\text{dim}(a)$
@@
SQIL Labels (Section B-F (1,1)(-1,1)
@@
DAC reward function: for VPACE-DAC and VP-DAC, although there are many options for reward functions that map $D$ to $\hat{R} \citeappghasemipourDivergenceMinimizationPerspective$2019, following [4, 12, 15], we set the reward to $\hat{R}_{\mathcal{T}}(s)=\log\left(D_{\mathcal{T}}(s)\right)-\log\left(1-D_{\mathcal{T}}(s)\right)$.
@@
Reward Scaling of .1: we use a reward scaling parameter of .1 for all implementations. Coupled with a discount rate $\gamma=0.99$ (common for much work in RL), this sets the expected minimum and maximum Q values for SQIL to $\frac{-.1}{1-\gamma}=-10$ and $\frac{.1}{1-\gamma}=10$.
@@ B-F1 Maintaining $Q^{\pi}_{\text{max}}$, $Q^{\pi}_{\text{min}}$ Estimates for Value Penalization
Our approach to value penalization requires maintaining estimates for or choosing $Q^{\pi}_{\text{max}}$ and $Q^{\pi}_{\text{min}}$. In both DAC and SQIL, the estimate of $Q^{\pi}_{\text{max}}$ comes from taking the mini-batch of data from $\mathcal{B}^{*}$, passing it through the Q function, taking the mean, and then using a median moving average filter to maintain an estimate. The $Q^{\pi}_{\text{max}}$, $Q^{\pi}_{\text{min}}$ num. filter points” value from Table IV refers to the size of this filter. We chose 50 and used it for all of our experiments. We set $Q^{\pi}_{\text{min}}$ to
@@
[…]